OpenStudy (anonymous):

y varies directly with x. When y = 12, x = 16. Find x when y = 24

OpenStudy (anonymous):

y=kx in direct relationships (k is the constant of variation)

OpenStudy (anonymous):

thus 12=k x 16 so k=12/16= 3/4

OpenStudy (anonymous):

finally 24=3/4x X and solve for X=32